# Geometrical Aspects Of Resources Distribution In Quantum Random Circuits

### v2

@@ 2 Quantum Random Circuits: distribution of resources
We start by analyzing a simple two-qubits system and the violation of CHSH inequalities [19]. For the state associated to each generated QRC, we compute its maximal degree of the violation. We consider from now on an experimental situation in which two dichotomic observables $A_{1}$, $A_{1}^{\prime}$, $A_{2}$ and $A_{2}^{\prime}$, with outputs $1$ and $-1$ each, can be performed on each system. Let
@@
where $E(A_{1}A_{2})$, for example, is the correlation function between $A_{1}$ and $A_{2}$, and so on with the primed variables. On the basis of the so called local realism (or local hidden variable model) it is possible to obtain the CHSH inequality [19]:
@@
Delving into the realm of quantum mechanics, our focus on dichotomic measurements allows us to confine our exploration to two-dimensional systems. In this case the operator inside the modulus of the left-hand side of Eq.2 becomes:
$O_{2}=\vec{\sigma}\cdot\vec{a}_{0}\otimes\vec{\sigma}\cdot\vec{b}_{0}-\vec{\sigma}\cdot\vec{a}_{1}\otimes\vec{\sigma}\cdot\vec{b}_{0}+\vec{\sigma}\cdot\vec{a}_{0}\otimes\vec{\sigma}\cdot\vec{b}_{1}+\vec{\sigma}\cdot\vec{a}_{1}\otimes\vec{\sigma}\cdot\vec{b}_{1},$ (3)
where $\vec{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$ with $\sigma_{x}$, $\sigma_{y}$ and $\sigma_{z}$ that are the Pauli operators and $\vec{a}_{0}$, $\vec{a}_{1}$, $\vec{b}_{0}$ and $\vec{b}_{1}$, are unit vectors representing the orientations of the polarizers of each pair. To each direction there is a pair of angles associated. For a given state $\rho$ we maximize the quantity
@@
among possible values of the angles. If the local realism is violated and consequentely non locality is present in a given state $\rho$ with regard to the chosen family of observables [^1] Notice that, in principle, there could exist a more general family of observables for which a non-locality inequality is violated. We are restricting here the analysis to the CHSH type observables (with variable angles)., there should exist a combination of angles for which $S_{2}(\rho)=V_{2}(\rho)>2$.
Now we turn to the study of systems with 2, 3, 4 and 5 qubits. When more than two particles are present, there exist different types of non-local states. First, one could start asking the question whether there exists a local hidden-variable model reproducing the correlations of the system under study. This problem was addressed by Mermin [14] (see also [20] and [21]). By requiring the condition of local hidden variable model, it is possible to derive a set of inequalities based on the so-called Mermin polynomials $M_{N}$ (see for example [20], section II-C). These polynomials are recursively defined in the following way
@@
where $M^{\prime}_{N-1}$ is obtained from $M_{N-1}$ by exchanging all the zero-indexed and and one-indexed $A$’s. The local realism limit is given by [20]:
@@
For $N=2$ we recover one of the equivalent ways of writing the CHSH inequalities up to factor $1/2$. The violation of the $N$-particle Mermin inequality implies that the correlations cannot be modeled using local realism. But, in a multi-partite system, restricting the analysis to the simple absence of local hidden-variable models might yield a narrow perspective on the possible correlations involved. For example, one could have a state of three particles in which the first two are maximally correlated, while there is no correlation with regard to the third one, as is the case for a quantum system of three particles prepared in the state
@@
The above state clearly displays non-local correlations between the first and second particles, but these correlations do not involve all the parties. In order to distinguish among the different types of correlations, G. Svelitchny and collaborators [15, 16] developed a set of inequalities which, if violated, will reveal the presence of genuine multipartite non-locality [^2] See also the discussion presented in Ref. [23].. The Svetlichny inequalities have also been experimentally confirmed (see for example [22]).
@@
$S_{3}$ $=E\left(A_{1}A_{2}A_{3}\right)+E\left(A_{1}A_{2}A_{3}^{\prime}\right)+E\left(A_{1}A_{2}^{\prime}A_{3}\right)+E\left(A_{1}^{\prime}A_{2}A_{3}\right)$ (9)
$-E\left(A_{1}^{\prime}A_{2}^{\prime}A_{3}^{\prime}\right)-E\left(A_{1}^{\prime}A_{2}^{\prime}A_{3}\right)-E\left(A_{1}^{\prime}A_{2}A_{3}^{\prime}\right)-E\left(A_{1}A_{2}^{\prime}A_{3}^{\prime}\right),$
@@
$O_{3}$ $=A_{1}\otimes A_{2}\otimes A_{3}+A_{1}\otimes A_{2}\otimes A_{3}^{\prime}+A_{1}\otimes A_{2}^{\prime}\otimes A_{3}+A_{1}^{\prime}\otimes A_{2}\otimes A_{3}$ (10)
$-A_{1}^{\prime}\otimes A_{2}^{\prime}\otimes A_{3}^{\prime}-A_{1}^{\prime}\otimes A_{2}^{\prime}\otimes A_{3}-A_{1}^{\prime}\otimes A_{2}\otimes A_{3}^{\prime}-A_{1}\otimes A_{2}^{\prime}\otimes A_{3}^{\prime}.$
@@
for all possible values of the polarizers. As in the two-qubits case, a violation of the above inequality implies that the state has genuine three-partite non-locality. Thus, given a state $\rho$, we look for the angles which maximize (11), and use that maximal value to quantify the eventual non-locality present in the state. We refer to the value obtained as violation level. A state with no non-locality with regard to the chosen set of observables, will display a violation level below the classical limit of 4. As is well known [15], for the three-qubits case, the maximum value of violation that can be reached with a quantum state is $4\sqrt{2}$. Thus, it is the maximum violation level reachable using quantum states.

### v3

@@ Acknowledgements
We start by analyzing a simple two-qubit system and the violation of CHSH inequalities [46]. For the state associated with each generated QRC, we compute its maximal degree of violation. We consider from now on an experimental situation in which two dichotomic observables $A_{1}$, $A_{1}^{\prime}$, $A_{2}$ and $A_{2}^{\prime}$, with outputs $1$ and 1-1 each, can be performed on each system. Let
@@
where $E(A_{1}A_{2})$, for example, is the correlation function between $A_{1}$ and $A_{2}$, and so on with the primed variables. Based on the so-called local realism (or local hidden variable model), it is possible to obtain the CHSH inequality [46]:
@@
Delving into the realm of quantum mechanics, our focus on dichotomic measurements allows us to confine our exploration to two-dimensional systems. In this case, the operator inside the modulus of the left-hand side of Eq. (2) becomes:
$\begin{split}O_{2}&=\vec{\sigma}\cdot\vec{a}_{0}\otimes\vec{\sigma}\cdot\vec{b}_{0}-\vec{\sigma}\cdot\vec{a}_{1}\otimes\vec{\sigma}\cdot\vec{b}_{0}+\vec{\sigma}\cdot\vec{a}_{0}\otimes\vec{\sigma}\cdot\vec{b}_{1}\\ &+\vec{\sigma}\cdot\vec{a}_{1}\otimes\vec{\sigma}\cdot\vec{b}_{1},\end{split}$ (3)
where $\vec{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$ with $\sigma_{x}$, $\sigma_{y}$ and $\sigma_{z}$ that are the Pauli operators and $\vec{a}_{0}$, $\vec{a}_{1}$, $\vec{b}_{0}$ and $\vec{b}_{1}$, are unit vectors representing the orientations of the polarizers of each pair. To each direction, there is a pair of angles associated. For a given state $\rho$ we maximize the quantity
@@
among possible values of the angles. If the local realism is violated and consequently non-locality is present in a given state $\rho$ with regard to the chosen family of observables  [^3] Notice that, in principle, there could exist a more general family of observables for which a non-locality inequality is violated. We are restricting here the analysis to the CHSH-type observables (with variable angles)., there should exist a combination of angles for which $S_{2}(\rho)=V_{2}(\rho)>2$.
Now we turn to the study of systems with $2$, 33, 44 and $5$ qubits. When more than two particles are present, there exist different types of non-local states. First, one could start asking the question of whether there exists a local hidden-variable model reproducing the correlations of the system under study. This problem was addressed by Mermin [16] (see also [47] and [48]). By requiring the condition of the local hidden variable model, it is possible to derive a set of inequalities based on the so-called Mermin polynomials $M_{N}$ (see for example [47], section II-C). These polynomials are recursively defined in the following way:
@@
where $M^{\prime}_{N-1}$ is obtained from $M_{N-1}$ by exchanging all the zero-indexed and one-indexed $A$’s. The local realism limit is given by [47]:
@@
For $N=2$ we recover one of the equivalent ways of writing the CHSH inequalities up to factor $1/2$. The violation of the $N$-particle Mermin inequality implies that the correlations cannot be modeled using local realism. However, in a multi-partite system, restricting the analysis to the simple absence of local hidden-variable models might yield a narrow perspective on the possible correlations involved. For example, one could have a state of three particles in which the first two are maximally correlated, while there is no correlation with regard to the third one, as is the case for a quantum system of three particles prepared in the state
@@
The above state clearly displays non-local correlations between the first and second particles, but these correlations do not involve all the parties. In order to distinguish among the different types of correlations, G. Svetlichny and collaborators [17, 18] developed a set of inequalities which, if violated, will reveal the presence of genuine multipartite non-locality (See also the discussion presented in Ref. [49]). The Svetlichny inequalities have also been experimentally confirmed (see for example [50]).
@@
$S_{3}=E\left(A_{1}A_{2}A_{3}\right)+E\left(A_{1}A_{2}A_{3}^{\prime}\right)+E\left(A_{1}A_{2}^{\prime}A_{3}\right)+E\left(A_{1}^{\prime}A_{2}A_{3}\right)-E\left(A_{1}^{\prime}A_{2}^{\prime}A_{3}^{\prime}\right)-E\left(A_{1}^{\prime}A_{2}^{\prime}A_{3}\right)-E\left(A_{1}^{\prime}A_{2}A_{3}^{\prime}\right)-E\left(A_{1}A_{2}^{\prime}A_{3}^{\prime}\right),$ (9)
@@
$O_{3}$ $=A_{1}\otimes A_{2}\otimes A_{3}+A_{1}\otimes A_{2}\otimes A_{3}^{\prime}+A_{1}\otimes A_{2}^{\prime}\otimes A_{3}$ (11)
$+A_{1}^{\prime}\otimes A_{2}\otimes A_{3}-A_{1}^{\prime}\otimes A_{2}^{\prime}\otimes A_{3}^{\prime}-A_{1}^{\prime}\otimes A_{2}^{\prime}\otimes A_{3}$
$-A_{1}^{\prime}\otimes A_{2}\otimes A_{3}^{\prime}-A_{1}\otimes A_{2}^{\prime}\otimes A_{3}^{\prime}.$
@@
for all possible values of the polarizers. As in the two-qubits case, a violation of the above inequality implies that the state has genuine three-partite non-locality. Thus, given a state $\rho$, we look for the angles which maximize (12) and use that maximal value to quantify the eventual non-locality present in the state. We refer to the value obtained as violation level. A state with no non-locality with regard to the chosen set of observables will display a violation level below the classical limit of 4. As is well known [17], for the three-qubit case, the maximum value of violation that can be reached with a quantum state is $4\sqrt{2}$. Thus, it is the maximum violation level reachable using quantum states.